# Electricity Market Bidding for Renewable Electrolyzer Plants: An Opportunity Cost Approach

### v2

@@ II-A1 Hydrogen production curve
The starting point of the derivation is the hydrogen production curve ${\rm h}(p^{\rm h})$ of the electrolyzer [18]. This curve expresses the amount of hydrogen ${\rm h}$ produced per hour at a given power consumption $p^{\rm h}$. Here, we leverage the empirical hydrogen production curve for an alkaline electrolyzer from [8], which is shown in Fig. 1(a). Due to low efficiencies at low and high loads, the hydrogen production curve is non-convex.
II-A2 Hydrogen revenue
@@
For an electrolyzer within a REP where the renewable generation capacity is larger than the electrolyzer capacity, i.e., $P^{\rm RES}>\bar{P}^{\rm h}$, the available renewable power may at times exceed the electrolyzer capacity. In these cases, the excess power cannot be used to further increase hydrogen production. The maximum hydrogen revenue is therefore equal to the revenue of maximum hydrogen production, ${\rm r}_{\rm h}(\bar{P}^{\rm h})$. To reflect this case, we formulate a piece-wise curve, defining the hydrogen revenue for any available (positive) power:
${\rm r}_{\rm h}(p^{\rm h})=\begin{cases}{\rm h}(p^{\rm h})\lambda^{\rm{h}},&p^{\rm h}\in[0,\bar{P}^{\rm h}],\\ {\rm h}(\bar{P}^{\rm h})\lambda^{\rm{h}},&p^{\rm h}>\bar{P}^{\rm h}.\end{cases}$ (2)
The first piece follows what is illustrated in Fig. 1(a), and the second piece is a constant equal to the revenue of maximum hydrogen production.
II-A3 Opportunity cost
For the REP to sell power rather than produce hydrogen, the revenue from selling power must exceed the revenue that the REP would get from withholding power for hydrogen production. The REP’s cost of selling a power $p^{\rm DA}$ is thus an opportunity cost, equal to the lost revenue of not producing hydrogen from the sold power.
To obtain the opportunity cost, we first derive the revenue the REP can achieve by withholding all the available renewable power $P^{\rm RES}$. We insert $P^{\rm RES}$ into (2) and obtain ${\rm r}_{\rm h}(P^{\rm RES})$, which for a given available renewable power is a constant. Following ${\rm r}_{\rm h}$’s piece-wise definition, ${\rm r}_{\rm h}(P^{\rm RES})$ is capped at the revenue of maximum hydrogen production if $P^{\rm RES}$ exceeds the electrolyzer capacity $\bar{P}^{\rm h}$.
Next, we look at the revenue achieved from hydrogen production as a function of power sold to the DA market. When selling a power $p^{\rm DA}$, the power available for hydrogen production is given by $p^{\rm h}=P^{\rm RES}-p^{\rm DA}$. Inserting into (2), we achieve ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$. Again, this revenue is capped at the revenue associated with maximum hydrogen production and takes this constant value when $P^{\rm RES}-p^{\rm DA}>\bar{P}^{\rm h}$.
The opportunity cost ${\rm c}_{\rm el}(p^{\rm DA})$ of selling a power $p^{\rm DA}$ is equal to the difference between the hydrogen revenue of withholding all available renewable power and the hydrogen revenue when selling a power $p^{\rm DA}$,
@@ II-A3 Opportunity cost
where $p^{\rm DA}$’s domain depends on the available renewable power relative to the electrolyzer capacity,
@@
Following from eq. (4), the REP can always export its full available renewable capacity, but can also bid to import power, i.e., have a negative $p^{\rm DA}$, if its available renewable production $P^{\rm RES}$ is lower than the electrolyzer capacity $\bar{P}^{\rm h}$. In reality, a negative $p^{\rm DA}$ represents a willingness-to-pay to import power, and not a cost. For simplicity, we still refer to the full curve as the (opportunity) cost curve.
The cost curve, plotted for the case where $P^{\rm RES}=\bar{P}^{\rm h}$ in Fig. 1(c), is a shifted reflection of the hydrogen revenue curve over both axes. First, as the cost curve, a function of $p^{\rm DA}$, is a constant ${\rm r}_{\rm h}(P^{\rm RES})$ subtracted by the revenue function ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$, its resulting curve is a vertical reflection of the hydrogen revenue curve, shifted along the y-axis with the value of ${\rm r}_{\rm h}(P^{\rm RES})$. Further, as the variable $p^{\rm DA}$ is negatively represented in the hydrogen revenue function, ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$, it is also a horizontal reflection of the hydrogen revenue curve, shifted along the x-axis with the constant $P^{\rm RES}$. This implies that as the hydrogen revenue curve is (approximately) concave, the REP’s cost curve is (approximately) convex.
In the cases where $P^{\rm RES}$ exceeds the electrolyzer capacity, ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$ is equal to the maximum hydrogen revenue ${\rm r}_{\rm h}(P^{\rm RES})$ for $p^{\rm DA}\in[0,P^{\rm RES}-\bar{P}^{\rm h}]$. For this range of sold power, the opportunity cost function thus evaluates to zero, indicating that there is no opportunity cost associated with selling renewable power that exceeds what the electrolyzer can consume. The cost curve then forms a piece-wise function, as follows from eq. (2).
The presented method allows for computing the cost curve of a stand-alone electrolyzer, by setting $P^{\rm RES}$ to zero.
@@
II-A4 Marginal opportunity cost
We find the REPs marginal cost of selling power ${\rm mc}_{\rm el}(p^{\rm DA})$ as the change in opportunity cost $d{\rm c}_{\rm el}$ over the change in sold power $dp^{\rm DA}$, i.e., the derivative of the opportunity cost,
@@ II-A4 Marginal opportunity cost
II-B Piecewise-linear approximation of the cost curve
As seen in Figs. 1(c),1(d), the REP (marginal) cost curve is non-convex. However, most electricity markets require participants to provide piecewise linear, convex cost curves, which corresponds to piecewise constant, non-decreasing marginal costs. We therefore derive an approximated cost curve.
Instead of directly approximating the opportunity cost curve in Fig. 1(c), we begin by approximating the hydrogen production curve in Fig. 1(a). As will become evident, this allows us to express our approximated opportunity cost curve directly as a function of the hydrogen price $\lambda^{\rm{h}}$, which allows for easier modification in case of hydrogen price changes. We approximate the hydrogen production curve as a concave piecewise linear function with decreasing slope:
@@ II-B Piecewise-linear approximation of the cost curve
where $A_{i}$ and $B_{i}$ represent the slope and constant corresponding to the $i^{th}$ linear piece, and $[\underline{P}_{i},\overline{P}_{i}]$ is the range of electrolyzer power consumption corresponding to piece $i$. The set $\mathcal{I}$ represents the set of pieces, and $\mathbf{1}_{(.)}$ is the indicator function which takes on a value of 1 when $p^{\rm{h}}\in[\underline{P}_{i},\overline{P}_{i}]$ (and zero otherwise). The approximation is chosen such that the slope is decreasing for all pieces, i.e., $A_{i}>A_{i+1}$.
From the approximated hydrogen production curve we derive the approximated opportunity cost for selling $p^{\rm DA}$:
@@
$=\lambda^{\rm h}\sum_{i\in\mathcal{I}}\left(A_{i}P^{\rm RES}+B_{i}\right)\mathbf{1}_{\left(P^{\rm{RES}}\in[\overline{P}_{i},\underline{P}_{i}]\right)}$ (7)
$-\lambda^{\rm h}\sum_{i\in\mathcal{I}}\left(A_{i}(P^{\rm RES}-p^{\rm DA})+B_{i}\right)\mathbf{1}_{\left((P^{\rm{RES}}-p^{\rm{DA}})\in[\overline{P}_{i},\underline{P}_{i}]\right)}.$
Since we approximated the hydrogen production curve as a concave, piecewise linear function, the resulting cost curve forms a convex, piecewise linear function, with increasing slopes. It can therefore be directly implemented into, for example, a DC OPF, without the need for binary variables.
The marginal opportunity cost is the derivative of the opportunity cost, which for the piece-wise linear curve is the slope of each piece with respect to $p^{\rm DA}$. The approximated marginal cost therefore becomes an increasing, step-wise function:
${\rm mc}_{\rm el}(p^{\rm{DA}})=\sum_{i\in\mathcal{I}}\lambda^{\rm h}A_{i}\mathbf{1}_{\left((P^{\rm{RES}}-p^{\rm{DA}})\in[\overline{P}_{i},\underline{P}_{i}]\right)}.$ (8)
@@
  • If there is excess renewable power production, $P^{\rm RES}>\bar{P}^{\rm h}$, the cost curves are shifted right to higher values of $p^{\rm DA}$. For the excess renewable generation, we define an initial piece with $\underline{P}_{1}=0,\overline{P}_{1}=P^{\rm{RES}}-\overline{P}^{\rm h}$, for which the cost is zero, i.e., $A_{1}=b_{1}=0$. This case is illustrated in the second column of graphs in Fig. 2.
@@
II-C Integrating Electrolyzers in the Electricity Market Clearing
The market-bidding REP submits its approximated cost curve, derived in the previous section, to the system operator and receives a generation schedule $p^{\rm DA}$ after the electricity market has cleared. The REP then uses the remaining power for hydrogen production (up to the full capacity of the electrolyzer), i.e., $p^{\rm h}=\min\left(P^{\rm RES}-p^{\rm DA},\bar{P}^{\rm h}\right)$.
@@ II-C Integrating Electrolyzers in the Electricity Market Clearing
where the variables $c_{n}^{\rm G}$, $p_{n}^{\rm G}$ and $\theta_{n}$ denote the total generation cost, power generation schedule, and voltage angle at node $n$, respectively. The objective function (9a) minimizes the total cost of generation. Constraint (9b) implements the piecewise linear, convex cost curves with parameters $\alpha_{i}$ and $\beta_{i}$, representing the slope and intercepts for each piece. Note that the inequality will always be binding (i.e., satisfied with equality) as the cost variables $c_{n}^{\rm G}$ are minimized. The reactance of each branch is given by ${X_{nm}}$, where $m\in\mathcal{K}_{n}$ notes all nodes $m$ connected to $n$. Constraint (9c) enforces the nodal power balance, while (9d) and (9e) ensure minimum and maximum limits on branch power flow and generation capacity. Finally, (9f) defines the reference node. The Locational Marginal Prices (LMPs) in the day-ahead stage are given by the dual variable associated with the power balance constraints (9c), denoted as $\lambda^{\rm DA}_{n}$. REPs are modeled as generators in this formulation, with the day-ahead schedule $p^{\rm DA}$ represented by a generator variable $p_{n}^{\rm G}$. Here, we use the slope and intercepts as derived in (II-B) to represent the REP. However, assuming no start-up costs, the approximated marginal cost curve is equivalent to the full cost curve and can also be implemented in a DC-OPF. Further, the marginal cost curve is important for understanding the REP’s position in the merit order curve.
@@
III Case study
This section analyzes the impact of connecting an electrolyzer to the grid, when co-located with a renewable power plant. We compare active market participation through the proposed bidding approach to the traditional approach of fixed electrolyzer consumption.
III-A Case Study Description
We define our case study based on the Reliability Test System—Grid Modernization Lab Consortium (RTS-GMLC) [19], an updated version of the IEEE RTS-96 with high shares of renewable power capacity. The RTS-GMLC system is divided into three regions, and comes with one year of data for hourly load and renewable production, as well as piece-wise cost data for all conventional generators. The load in the system is inflexible. We refer to the RTS-GMLC system without any modifications, i.e., without any electrolyzer capacity, as the base system. The production capacities and demand range of each region in the system are plotted in Fig. 3. Note that the majority of the renewable capacity in the system is located in region 3.

### v3

@@ III-B1 Hydrogen production curve
This curve expresses the amount of hydrogen ${\rm h}$ produced per hour at a given power consumption $p^{\rm h}$, and does not have a closed form expression. Here, we leverage the empirical hydrogen production curve for an alkaline electrolyzer from [10], which is shown in Fig. 2(a). Due to low efficiencies at low and high loads, the hydrogen production curve is non-convex.
III-B2 Hydrogen revenue
@@
For an electrolyzer within a REP where the renewable generation capacity is larger than the electrolyzer capacity, i.e., $P^{\rm RES}>\bar{P}^{\rm h}$, the available renewable power may at times exceed the electrolyzer capacity. In these cases, the excess power cannot be used to further increase hydrogen production. The maximum hydrogen revenue is therefore equal to the revenue of maximum hydrogen production, ${\rm r}_{\rm h}(\bar{P}^{\rm h})$. To reflect this case, we formulate a piecewise curve, defining the hydrogen revenue for any available (positive) power $p$:
${\rm r}_{\rm h}(p)=\begin{cases}{\rm h}(p)\lambda^{\rm{h}},&p\in[0,\bar{P}^{\rm h}],\\ {\rm h}(\bar{P}^{\rm h})\lambda^{\rm{h}},&p>\bar{P}^{\rm h}.\end{cases}$ (3)
The first piece of eq. (3) follows what is illustrated in Fig. 2(a), and the second piece is a constant equal to the revenue of maximum hydrogen production.
III-B3 Opportunity cost
For the REP to sell power instead of producing hydrogen, the revenue from selling power must exceed the revenue that the REP would earn from withholding power for hydrogen production. The REP’s cost of selling a power quantity $p^{\rm DA}$ is thus an opportunity cost, equal to the lost revenue of not producing hydrogen from the sold power. The sold power $p^{\rm DA}$ relates to the power consumed by the electrolyzer $p^{\rm h}$ through the internal power balance of the REP:
$P^{\rm RES}\geq p^{\rm DA}+p^{\rm h}.$ (4)
The opportunity cost ${\rm c}_{\rm el}(p^{\rm DA})$ of selling a power quantity $p^{\rm DA}$ is equal to the difference between the hydrogen revenue of withholding all available renewable power and the hydrogen revenue when selling a power $p^{\rm DA}$:
@@ III-B3 Opportunity cost
where ${\rm r}_{\rm h}(.)$ is the hydrogen revenue function defined in eq. (3). Here, ${\rm r}_{\rm h}(P^{\rm RES})$ represents the revenue from using all available renewable power for hydrogen production, while ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$ represents the revenue when a power quantity $p^{\rm DA}$ is sold to the electricity market, since this amount can no longer be used for hydrogen production. The range of possible $p^{\rm DA}$ depends on the available renewable power relative to the electrolyzer capacity. The REP can always export power $p^{\rm DA}$ up until its available renewable capacity $P^{\rm RES}$. However, if its available renewable production $P^{\rm RES}$ is less than or equal to the electrolyzer capacity $\bar{P}^{\rm h}$, i.e., $P^{\rm RES}\leq\bar{P}^{\rm h}$, the REP is also allowed to bid to import power. This is expressed mathematically as:
@@
In reality, a negative $p^{\rm DA}$ represents a willingness-to-pay to import power, and not a cost. For simplicity, we still refer to the full curve as the (opportunity) cost curve.
The cost curve, plotted for the case where $P^{\rm RES}=\bar{P}^{\rm h}$ in Fig. 2(c), is a shifted reflection of the hydrogen revenue curve ${\rm r}_{\rm h}$(.) over both axes. The vertical reflection arises from the fact that the hydrogen revenue curve is subtracted from a constant ${\rm r}_{\rm h}(P^{\rm RES})$ in the cost curve. The horizontal reflection comes from the fact that the variable $p^{\rm DA}$ appears negatively in the hydrogen revenue function ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$. This implies that as the hydrogen revenue curve is (approximately) concave, the REP’s cost curve is (approximately) convex.
In the cases where $P^{\rm RES}$ exceeds the electrolyzer capacity, ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$ is equal to the maximum hydrogen revenue ${\rm r}_{\rm h}(P^{\rm RES})$ for $p^{\rm DA}\in[0,P^{\rm RES}-\bar{P}^{\rm h}]$. For this range of sold power, the opportunity cost function thus evaluates to zero, indicating that there is no opportunity cost associated with selling renewable power that exceeds what the electrolyzer can consume. The cost curve then forms a piecewise (non-linear) function. The presented method allows for computing the cost curve of a stand-alone electrolyzer, by setting $P^{\rm RES}$ to zero.
@@
III-B4 Marginal opportunity cost
We find the REP’s marginal cost of selling power ${\rm mc}_{\rm el}(p^{\rm DA})$ as the change in opportunity cost $d{\rm c}_{\rm el}$ over the change in sold power $dp^{\rm DA}$, i.e., the derivative of the opportunity cost,
@@ III-B4 Marginal opportunity cost
III-C Piecewise-linear approximation of the cost curve
As illustrated in Figs. 2(c) and 2(d), the REP’s (marginal) cost curve is non-convex. However, as argued in Section II, many electricity markets require participants to provide piecewise linear, convex cost curves, which corresponds to piecewise constant, non-decreasing marginal costs. We therefore derive an approximated cost curve.
Instead of directly approximating the opportunity cost curve in Fig. 2(c), we begin by approximating the hydrogen production curve in Fig. 2(a). As will become evident, this allows us to express our approximated opportunity cost curve directly as a function of the hydrogen price $\lambda^{\rm{h}}$, making it easier to adjust in response to changes in the hydrogen price. We approximate the hydrogen production curve as a concave, piecewise linear function with a decreasing slope, following the approach in [10], as described in Section III-A:
@@ III-C Piecewise-linear approximation of the cost curve
where $A_{i}$ and $B_{i}$ represent the slope and constant corresponding to the $i^{\rm th}$ linear piece, and $[\underline{P}_{i},\overline{P}_{i}]$ is the range of electrolyzer power consumption corresponding to piece $i$. The set $\mathcal{I}$ represents the set of pieces =$\mathcal{I}=\{1,2,\ldots,I\}$, and $\mathbf{1}_{(.)}$ is the indicator function which takes on a value of 1 when $p^{\rm{h}}\in[\underline{P}_{i},\overline{P}_{i}]$ (and zero otherwise). The approximation is chosen such that the slope is decreasing for all pieces, i.e., $A_{i}>A_{i+1}$.
From the approximated hydrogen production curve, we derive the approximated opportunity cost for selling $p^{\rm DA}$:
@@
$=\lambda^{\rm h}\sum_{i\in\mathcal{I}}\left(A_{i}P^{\rm RES}+B_{i}\right)\mathbf{1}_{\left(P^{\rm{RES}}\in[\underline{P}_{i},\overline{P}_{i}]\right)}$ (9)
$-\lambda^{\rm h}\sum_{i\in\mathcal{I}}\left(A_{i}(P^{\rm RES}-p^{\rm DA})+B_{i}\right)\mathbf{1}_{\left((P^{\rm{RES}}-p^{\rm{DA}})\in[\underline{P}_{i},\overline{P}_{i}]\right)}.$
The marginal opportunity cost is the derivative of the opportunity cost, which for the piecewise linear curve is the slope of each piece with respect to $p^{\rm DA}$. The approximated marginal cost therefore becomes an increasing, step-wise function:
${\rm mc}_{\rm el}(p^{\rm{DA}})=\sum_{i\in\mathcal{I}}\lambda^{\rm h}A_{i}\mathbf{1}_{\left((P^{\rm{RES}}-p^{\rm{DA}})\in[\underline{P}_{i},\overline{P}_{i}]\right)}.$ (10)
Since we approximated the hydrogen production curve as a concave, piecewise linear function, the resulting cost curve forms a convex, piecewise linear function, with increasing slopes. It can therefore be directly implemented into, for example, a DC OPF, without the need for binary variables. The slope and intercepts of the resulting piecewise linear cost curve, which form a convex hull, can be expressed through the slope and intercepts of the approximated hydrogen production curve. We achieve a piecewise linear approximation of the REP’s cost function ${\rm c_{el}}(p^{\rm DA})$,
${\rm c_{el}}(p^{\rm DA})=\alpha_{i}p^{\rm DA}+\beta_{i},\quad p^{\rm DA}\in[P^{\rm RES}-\overline{P}_{i},\;P^{\rm RES}-\underline{P}_{i}],$ (11)
for $i\in\mathcal{I}$ with slopes $\alpha_{i}$ and intercepts $\beta_{i}$,
$\alpha_{i}=\lambda^{\rm{h}}A_{i},$ (12a)
$\beta_{i}=\lambda^{\rm{h}}\bigl{(}(A_{j}-A_{i})P^{\rm RES}+(B_{j}-B_{i})\bigr{)},$ (12b)
where $j$ is the index such that $P^{\rm RES}\in[\underline{P}_{j},\overline{P}_{j}]$. Note that the slope in eq. (12a) corresponds to the marginal cost defined in eq. (10). In some cases, the RES production may exceed the electrolyzer capacity. To capture this, we define an additional piece on the cost curve. We expand the set of pieces on the approximated cost curve so that =$\mathcal{I}=\{1,2,\ldots,I,I+1\}$ , where $I$ is the number of pieces on the hydrogen production curve. Further, we define that $A_{I+1}=0$, $B_{I+1}=A_{I}\overline{P}_{I}+B_{I}$, $\underline{P}_{I+1}=\bar{P}^{\rm h}$, and $\overline{P}_{I+1}=P^{\rm RES}$.
The full derivation of the steps from the original formulation of the REP’s cost curve in eq. (5) to the piecewise expressions in eqs. (11) and (12), as well as the definition of the additional piece $I+1$ is provided in Appendix -A.
@@
  • If there is excess renewable power production, $P^{\rm RES}>\bar{P}^{\rm h}$, the cost curves are shifted right to higher values of $p^{\rm DA}$. For the excess renewable generation, the additional piece defined evaluates to zero, as there is no opportunity cost of hydrogen production associated. This case is illustrated in the second column of graphs in Fig. 3.
@@
III-D Integrating Electrolyzers in the Electricity Market Clearing
The market-bidding REP submits its approximated cost curve, derived in the previous section, to the market operator and receives a generation schedule $p^{\rm DA}$ after the electricity market has cleared. The REP then uses the remaining power for hydrogen production (up to the full capacity of the electrolyzer), i.e., $p^{\rm h}=\min\left(P^{\rm RES}-p^{\rm DA},\bar{P}^{\rm h}\right)$.
@@ III-D Integrating Electrolyzers in the Electricity Market Clearing
where the variables $c_{n}^{\rm G}$, $p_{n}^{\rm G}$ and $\theta_{n}$ denote the total generation cost, power generation schedule, and voltage angle at node $n$, respectively. The objective function (13a) minimizes the total cost of generation. Constraint (13b) implements the piecewise linear, convex cost curves with parameters $\alpha_{i}$ and $\beta_{i}$, representing the slope and intercepts for each piece. Note that the inequality will always be binding (i.e., satisfied with equality) as the cost variables $c_{n}^{\rm G}$ are minimized. The reactance of each branch is given by ${X_{nm}}$, where $m\in\mathcal{K}_{n}$ notes all nodes $m$ connected to $n$. Constraint (13c) enforces the nodal power balance, while (13d) and (13e) ensure minimum and maximum limits on branch power flow and generation capacity. Finally, (13f) defines the reference node. The Locational Marginal Prices (LMPs) in the day-ahead stage are given by the dual variable associated with the power balance constraints (13c), denoted as $\lambda^{\rm DA}_{n}$. REPs are modeled as generators in this formulation, with the day-ahead schedule $p^{\rm DA}$ represented by a generator variable $p_{n}^{\rm G}$. Here, we use the slope and intercepts as derived in (III-C) to represent the REP. However, assuming no start-up costs, the approximated marginal cost curve is equivalent to the full cost curve and can also be implemented in a DC-OPF. Further, the marginal cost curve is important for understanding the REP’s position in the merit order curve.
@@
[…]
Figure 4: Schematic of an REP actively participating in an electricity market (here, a day-ahead market). An exogenous, fixed hydrogen price $\lambda^{\rm h}$, combined with known local renewable power production $P^{\rm RES}_{t}$ at hour $t$ and hydrogen production curve ${\rm h}(p^{\mathrm{h}})$, defines the REP’s bid curve for that hour. This curve is characterized by piecewise slopes $\alpha_{it}$ and intercepts $\beta_{it}$, where $i$ indexes the segments. The bid is submitted to the electricity market, together with the cost curves of other generators and the inflexible demand at the time $P^{\rm D}_{nt}$, which returns the clearing prices $\lambda^{\rm DA*}_{nt}$ and production schedules $p^{\rm G*}_{nt}$.
Fig. 4 summarizes how the REP is modeled as a market participant in this work. The REP receives an exogenous hydrogen price, $\lambda^{\rm h}$. This price, together with the (forecasted) internal RES production at time $t$, $P^{\rm RES}_{t}$, and the electrolyzer’s hydrogen production curve ${\rm h}(p^{\rm h})$, is used to derive the REP’s piecewise linear bid curve. This bid curve is given by slopes $\alpha_{it}$ and intercepts $\beta_{it}$ for piece $i$ on the curve. The bid curve is submitted to the electricity market, here modeled as a DC OPF (13), which outputs the cleared schedules $p^{\rm G*}_{nt}$ and prices $\lambda^{\rm DA*}_{nt}$ of all generators and nodes in the system.
IV Case study
This section analyzes the impact of connecting an electrolyzer to the grid, when co-located with a renewable power plant, comparing active market participation through the proposed bidding approach to a traditional fixed electrolyzer consumption.
IV-A Case Study Description
We define our case study based on the Reliability Test System—Grid Modernization Lab Consortium (RTS-GMLC) [37], an updated version of the IEEE RTS-96 with high shares of renewable power capacity. The RTS-GMLC system is divided into three regions, and comes with one year of data for hourly load and renewable production, as well as piecewise cost data for all conventional generators. The load in the system is inflexible. We refer to the RTS-GMLC system without any modifications, i.e., without any electrolyzer capacity, as the base system. The production capacities and demand range of each region in the system are plotted in Fig. 5. Note that the majority of the renewable capacity in the system is located in region 3.